# 3VL: Using Trees to Improve Vision-Language Models’ Interpretability

### v1

@@ V Experiments
We start by evaluating 3VL on the VL-Checklist. In the VL-Checklist [63] dataset we have a set of images and a pair of positive and negative captions for each image. The model is tested for image-text retrieval. The goal of this benchmark is to evaluate VLMs’ compositional understanding of objects, attributes and relations. Each negative caption is created by replacing one word in the positive sentence with respect to objects, relations or attributes.
We evaluate 3VL performance with and without Token Removal. Table I presents the results on the VL-Checklist VG datasets [63], namely, VG Relation, Object and Attribute. Note that the VG datasets are the most challenging ones in the VL-Checklist benchmark. In the experiments, we compare our approach to openAI VIT-B/32 CLIP [48], NegCLIP [98], RB+LLM Negs [1]. The evaluation of the models’ performance involved measuring the Top-1 accuracy metric. Note the improvement in 3VL compared to the CLIP baseline and that our approach improves over the other techniques in most cases.
@@
Table IX summarizes the results for the group categories of VL-Checklist. Notice that 3VL is significantly better than the other techniques on the relational and object categories. This may be attributed to the tree structure we introduce during training that helps in such type of queries. Yet, in attributes the improvement is better than NegCLIP but less than [1]. We conjecture that creating shallow rule-based hard negatives is more efficient for teaching attributes concepts like color and material while our tree structure is more beneficial for complex relational knowledge.
In addition, we report the results of 21 zero-shot tasks (ZS-Tasks). This includes Imagenet and the 20 tasks of ELEVATER [99]. This test checks the zero-shot image classification performance of our model and it shows that there is only little forgetting in our learning process. 3VL is better in this respect than NegCLIP but slightly worse than [1]. We also compare to CLIP and CLIP+LoRA finetuned on MS-COCO with contrastive loss (without tree augmented loss) as an ablation of the effects of finetuning on MSCOCO.
@@
The VSR [65] benchmark evaluates VLMs’ spatial reasoning using image-to-text retrieval. Each image contains a visible spatial relation, and for each image we have a pair of captions with a description of this visible spatial relation and textual “True” and “False” labels concatenated at the end (e.g. “A is in front of B (False)”, “A is in front of B (True)”). The model needs to retrieve the caption with the correct label.
Table III presents the results of the VSR [65] dataset with openAI CLIP VIT-B/32, NegCLIP [98], RB+LLM Negs [1], and 3VL. Although this task is very hard for all models (probably somewhat because of the unusual caption + label prompt style) we can still see an advantage for 3VL.
Figure 5: A comparison of vanilla CLIP (right) and 3VL (left) interpretability for a randomly selected sample from the VL-Checklist VG attribute action dataset, where both models are mistaken. At the top we present the HilaCAM [22] relevancy map visualizations (red areas important, purple not important) for the positive text (“airborne player”) and negative text (“standing player”) and the DiRe relevancy map. On the second row from left to right we can see the Token Removal process. Starting from the original image we gradually remove the least significant tokens according to the DiRe relevancy map.
@@
Image-Text retrieval. In Tables IV and V we report results of image to text (I2T) and text to image (T2I) retrieval on COCO and FLICKR datasets.
Image Segmentation. Table VI reports segmentation results using CLIPSeg [100] with vanilla CLIP and with 3VL on COCO and PhraseCut [101] datasets.
@@ V-D Interpretability Quantitative Assessment
To assess the interpretability of our method, we follow the negative perturbation (Token Removal) test in [22]. First a relevancy map is extracted from the image-text pair. Then, we remove image tokens in order of increasing importance (from lowest to highest importance according to the relevancy map) and use the image with removed tokens as input to the image encoder and calculate the cosine similarity to both the positive text and negative text. As we have two texts, positive and negative, for HilaCAM [22] we calculate one relevancy map for the positive text and one for the negative text. Then, we remove tokens according to the relevancy map created by each of the texts and get two subsets of tokens which we calculate cosine similarity for. For Anchor and DiRe we have only one relevancy map and we use it for Token Removal. The prediction in all cases is the text with the highest similarity score to the input image after Token Removal.
@@
V-D1 Interpretability User Study
@@
To further assess the interpretability of 3VL we conducted a user study. In the user study we have compared 3VL to vanilla CLIP using relevancy maps generated from image-text pairs with HilaCAM, and we have compared relevancy maps generated by HilaCAM to relevancy maps generated by Anchor and DiRe. We have compared the different relevancy maps by letting users decide which relevancy map emphasizes a caption in an image better. We randomly sampled 25 images from the VL-Checklist VG dataset. For each image we presented the original image on the right and the relevancy maps (either HilaCAM relevancy maps for 3VL and vanilla CLIP, or the relevancy maps of HilaCAM, Anchor and DiRe) in random order. A total of 37 respondents participated in the study and were asked “which image better emphasizes the caption…” (a), (b), (c) or neither. We have found 3VL maps to be more explainable than vanilla CLIP. 51.67% of user choices were 3VL, 27.4% vanilla CLIP and 20.93% neither. Moreover, in 78.94% of the questions 3VL got the most votes. We have also found DiRe to be more explainable than Anchor, which is more explainable than HilaCAM. 57.6% of user choices were DiRe, 19.2% Anchor, 11.2% HilaCAM and 12% neither. Moreover, in 83.33% of the questions DiRe got the most votes and in 16.67% of the questions Anchor got the most votes. HilaCAM did not get the most votes in any of the questions. In Figure  4 we present some examples from the user study.
V-E Interpretability Qualitative assessment
Figure 5 visualizes the strong interpretability of 3VL with HilaCAM and DiRe. We compare the visualizations of vanilla CLIP and 3VL with HilaCAM [22] and 3VL on a random sample from the VL-Checklist dataset, where both CLIP and 3VL are mistaken. Note that from the HilaCAM visualizations it is hard to understand the source of the error. With 3VL the visualization of the positive text emphasizes an airborne player and the visualization of the negative text emphasizes both an airborne player and a standing player. With Vanilla CLIP the positive and negative visualizations emphasize different parts of airborne players. Yet both models mistakenly predict “standing player”. When adding DiRe visualization we understand that the model should have focused more on the bottom part of the airborne player in addition to the upper part to understand that he is airborne. The Token Removal emphasizes this further. Figure 6 visualizes the strong interpretability of 3VL with HilaCAM and Anchor. 3VL attends to all lights in the photo, whereas Vanilla CLIP attends more to lights above seats. After Token removal, both attend to lights above seats.
V-F Using 3VL for Understanding VLMs’ Failures
Using our caption tree generation method we are able to find exact words and part of speech tags that the model fails to recognize. To offer extra insight regarding model’s decisions, we expand the caption tree at the level in which the model failed, with extra words generated to be similar (co-hyponyms and synonyms) to the chosen negative word and to the positive word and check probability for all these words. We also add relative relevancy visualizations for these words at failure points. Figure 7 demonstrates how CLIP fails to understand multiple words phrases and considers each single word on its own. Figure 8 shows the relative relevancy maps of the 4 most probable captions according to CLIP, highlighting different parts of the image for each(e.g. the wooden bench and the metal bench rail).
In Table VII we show statistics of failures per part-of-speech tag on MSCOCO test set with pretrained CLIP, ARO NegCLIP [98] and 3VL. Note the big improvement in some of CLIP’s biggest weaknesses, verbs and adpositions understanding, with more than 50% improvement on verbs and 46% improvement on adpositions.
To gain a better understanding of the model’s success and failures, we analyze the more frequent words failures and find some strong biases in vanilla CLIP. We show in Table VIII some CLIP biases and how it changes with 3VL.
@@ V-G Ablation Study
We conduct an ablation study to assess the impact of different factors on the effectiveness of our proposed tree-based training technique. We examine the effects of different loss functions, different tree structures and negatives generation. We examine the regularization effects of limiting our caption tree to a maximum depth or limiting the number of maximum negatives generated in each tree.
Different loss functions ablation. In Table IX we show the impact of different loss functions on the performance of our model. Note that the addition of LoRA degrades a bit the performance on VL-checklist but reduces the forgetting effect (exhibited in the performance on ELEVATER [99]). Note that the addition of tree loss leads to a significant improvement over just using contrastive learning.
TABLE IX: Top-1 accuracy on VL-Checklist(VLC) Relation(Rel) average, Attribute(Attr) average, Object(Obj) average, and average zero-shot(ZS) on ELEVATER 21 datasets. We test on openAI CLIP VIT-B/32, NegCLIP [98], RB+LLM Negs [1], and 3VL. As an ablation we also compare to CLIP and CLIP+LoRA that were trained without tree-based loss.
@@
      “LLM prompt” - we generate co-hyponyms, antonyms and random adjectives or adpositions with a prompt to FLAN-T5 (e.g., “find an opposite for the word: <>absent<>< >).
@@
In Table X we show Top-1 accuracy results on VL-Checklist with four such tree variants. (i) “Basic WN” - “Basic” tree structure with WordNet negatives generation, (ii) “Incremental WN” - “Incremental” tree structure with WordNet negatives generation, (iii) “WN+LLM prompt+mask” - “Incremental” tree structure with “LLM prompt” for opposites generation, T5 mask completion for words with no opposite and WordNet replacement if T5 mask completion generates a word with similar meaning. (iv) “WN+LLM prompt (3VL)” - “Incremental” tree structure with “LLM prompt” for opposites generation and WordNet replacement if no opposite exists.
@@
For each such variant, we report the highest result attained by hyperparameter sweep on the validation set of MSCOCO. We found that LLM mask completion oftentimes generated positive caption alternatives instead of negative captions and we did not use it in the final model. Best results were achieved using (iv) “WN+LLM prompt” and therefore, we used it in the final 3VL model.
@@
TABLE X: Top-1 accuracy on VL-Checklist for different caption trees. (i) “Basic WN” - Basic tree structure with WordNet replacement, (ii) “Incremental WN” - Incremental tree structure with WordNet replacement, (iii) “WN+LLM prompt+mask” - Incremental tree, FLAN-T5 prompt for opposites, T5 mask completion if opposite doesn’t exist and WordNet replacement if mask completion did not generate a word with a different meaning. (iv) (“WN+LLM prompt (3VL)”) - Incremental tree, FLAN-T5 prompt for opposites, and WordNet replacement if the opposite does not exist.
@@
Tree Method Att color Att material Att size Att action Att state Rel action Rel spatial Avg Basic WN 70.54 75.79 69.72 75.95 70.46 81.81 73.94 74.03 Incremental WN 74.41 79.23 70.13 78.12 71.63 83.81 75.70 76.15 WN+LLM prompt+mask 72.31 77.53 69.98 76.69 69.39 83.43 76.87 75.17 WN+LLM prompt (3VL) 75.29 82.63 70.90 81.10 75.03 81.69 81.14 78.25
@@ V-H Maximum tree depth ablation
In Table XI we report Top-1 accuracy on thr VL-Checklist when constraining the 3VL tree to maximum depth. We note that we don’t limit here the number of generated negatives per caption, just the structure of the tree. So, for example, a tree of depth 1 will contain the same number of negatives as in the full caption tree, just in a single level. We can see that even though the number of negatives remains the same, the increase in tree depth contributes to improved performance.

### v2

@@ V Experiments
We start by evaluating 3VL on the VL-Checklist [64] benchmark.
We evaluate 3VL performance with and without Token Removal. Table I presents the results on the VL-Checklist VG datasets [64], namely, VG Relation, Object, and Attribute. Note that the VG datasets are the most challenging ones in the VL-Checklist benchmark. In the experiments, we compare our approach to openAI VIT-B/32 CLIP [49], NegCLIP [101], RB+LLM Negs [1]. The evaluation of the models’ performance involved measuring the Top-1 accuracy metric. Note the improvement in 3VL compared to the CLIP baseline and that our approach improves over the other techniques in most cases.
@@
Table III summarizes the results for the group categories of VL-Checklist. Notice that 3VL is significantly better than the other techniques on the relational and object categories. This may be attributed to the tree structure we introduce during training that helps in such type of queries. Yet, in attributes, the improvement is better than NegCLIP but less than [1]. We conjecture that creating shallow rule-based hard negatives is more efficient for teaching attributes concepts like color and material while our tree structure is more beneficial for complex relational knowledge.
In addition, we report the results of 21 zero-shot tasks (ZS-Tasks). This includes Imagenet and the 20 image classification datasets of ELEVATER [102]. This test checks the zero-shot image classification performance of our model and it shows that there is only little forgetting in our learning process. 3VL is better in this respect than NegCLIP but slightly worse than [1]. We also compare to CLIP and CLIP+LoRA finetuned on MS-COCO with contrastive loss (without tree-augmented loss) as an ablation of the effects of finetuning on MSCOCO.
@@
The VSR [66] benchmark evaluates VLMs’ spatial reasoning using image-to-text retrieval. Each image contains a visible spatial relation, and for each image, we have a pair of captions with a description of this visible spatial relation and textual “True” and “False” labels concatenated at the end (e.g. “A is in front of B (False)”, “A is in front of B (True)”). The model needs to retrieve the caption with the correct label.
Table IV presents the results of the VSR [66] dataset with openAI CLIP VIT-B/32, NegCLIP [101], RB+LLM Negs [1], and 3VL. Although this task is very hard for all models (probably somewhat because of the unusual caption + label prompt style) we can still see an advantage for 3VL.
@@
Image-Text retrieval. In Tables V and VI we report results of image to text (I2T) and text to image (T2I) retrieval on COCO and FLICKR datasets.
Image Segmentation. Table VII reports segmentation results using CLIPSeg [103] with vanilla CLIP and with 3VL on COCO and PhraseCut [104] datasets.
Figure 6: A comparison of vanilla CLIP (right) and 3VL (left) interpretability for a randomly selected sample from the VL-Checklist VG attribute action dataset, where both models are mistaken. At the top we present the HilaCAM [22] relevancy map visualizations (red areas important, purple not important) for the positive text (“airborne player”) and negative text (“standing player”) and the DiRe relevancy map. On the second row from left to right, we can see the Token Removal process. Starting from the original image we gradually remove the least significant tokens according to the DiRe relevancy map.
@@ V-D Interpretability Quantitative Assessment
Assessing Interpretability via Token Removal. To assess the interpretability of our method, we follow the negative perturbation (Token Removal) test in [22]. First, a relevancy map is extracted from the image-text pair. Then, we remove image tokens in order of increasing importance (from lowest to highest importance according to the relevancy map) and use the image with removed tokens as input to the image encoder and calculate the cosine similarity to both the positive text and negative text. As we have two texts, positive and negative, for HilaCAM [22] we calculate one relevancy map for the positive text and one for the negative text. Then, we remove tokens according to the relevancy map created by each of the texts and get two subsets of tokens which we calculate cosine similarity for. For Anchor and DiRe we have only one relevancy map and we use it for Token Removal. The prediction in all cases is the text with the highest similarity score to the input image after Token Removal.
@@
The Token Removal evaluation shows that removal of less relevant tokens by our methods “Anchor” and DiRe” leads to greater performance improvement than token removal using other methods’ relevancy maps. Thus, we find that our generated relevancy maps are more “accurate” than others.
@@
Assessing Interpretability via User Study. To further assess the interpretability of 3VL we conducted a user study (for which we got the necessary IRB Approval). In the user study we have compared 3VL to vanilla CLIP using relevancy maps generated from image-text pairs with HilaCAM, and we have compared relevancy maps generated by HilaCAM to relevancy maps generated by Anchor and DiRe. We have compared the different relevancy maps by letting users decide which relevancy map emphasizes a caption in an image better. We randomly sampled 25 images from the VL-Checklist VG dataset. For each image we presented the original image on the right and the relevancy maps (either HilaCAM relevancy maps for 3VL and vanilla CLIP, or the relevancy maps of HilaCAM, Anchor and DiRe) in random order. A total of 37 respondents participated in the study and were asked “which image better emphasizes the caption…” (a), (b), (c) or neither. We have found 3VL maps to be more explainable than vanilla CLIP. 51.67% of user choices were 3VL, 27.4% vanilla CLIP and 20.93% neither. Moreover, in 78.94% of the questions 3VL got the most votes. We have also found DiRe to be more explainable than Anchor, which is more explainable than HilaCAM. 57.6% of user choices were DiRe, 19.2% Anchor, 11.2% HilaCAM and 12% neither. Moreover, in 83.33% of the questions DiRe got the most votes and in 16.67% of the questions Anchor got the most votes. HilaCAM did not get the most votes in any of the questions. Some screenshots from the user study appear in the appendix in Figure  10.
The user study shows that users have a preference for our relevancy maps. This may imply that our relevancy maps align more closely with human interpretation by highlighting important image parts. However, it is possible that while our relevancy maps are more visually appealing to users, they do not necessarily explain the model’s decision better. We leave a further exploration of the benefit of our relevancy map approach to a future research.
@@
V-E Interpretability Qualitative assessment
Figure 6 visualizes the strong interpretability of 3VL with HilaCAM and DiRe. We compare the visualizations of vanilla CLIP and 3VL with HilaCAM [22] and 3VL on a random sample from the VL-Checklist dataset, where both CLIP and 3VL are mistaken. Note that from the HilaCAM visualizations, it is hard to understand the source of the error. With 3VL the visualization of the positive text emphasizes an airborne player and the visualization of the negative text emphasizes both an airborne player and a standing player. With Vanilla CLIP the positive and negative visualizations emphasize different parts of airborne players. Yet both models mistakenly predict “standing player”. When adding DiRe visualization we understand that the model should have focused more on the bottom part of the airborne player in addition to the upper part to understand that he is airborne. The Token Removal emphasizes this further. Figure 7 visualizes the strong interpretability of 3VL with HilaCAM and Anchor. 3VL attends to all lights in the photo, whereas Vanilla CLIP attends more to lights above seats. After Token removal, both attend to lights above seats.
V-F Using 3VL for Understanding VLMs’ Failures
Using our caption tree generation method we are able to find exact words and part of speech tags that the model fails to recognize. To offer extra insight regarding the model’s decisions, we expand the caption tree at the level in which the model failed, with extra words generated to be similar (co-hyponyms and synonyms) to the chosen negative word and to the positive word and check the probability for all these words. We also add relative relevancy visualizations for these words at failure points. Figure 8 demonstrates how CLIP fails to understand multiple-word phrases and considers each single word on its own. Figure 9 shows the relative relevancy maps of the 4 most probable captions according to CLIP, highlighting different parts of the image for each(e.g. the wooden bench and the metal bench rail).
TABLE VIII: Number of failures on COCO testset per part of speech tag. In addition to the absolute number of failures, for NegCLIP and 3VL we present the % improvement over CLIP. For vanilla CLIP, NegCLIP [101] and 3VL. 3VL improves by more than 50% on verbs(VERB) and by 46% on adpositions(ADP).
Model NOUN ADP VERB ADJ CLIP 3104 2927 1612 913 NegCLIP 2743 / 11.63% 1989 / 32.04% 900 / 44.16% 695 / 23.87% 3VL 1976 / 36.34% 1575 / 46.19% 780 / 51.61% 612 / 32.96%
In Table VIII we show statistics of failures per part-of-speech tag on the MSCOCO test set with pretrained CLIP, ARO NegCLIP [101] and 3VL. Note the big improvement in some of CLIP’s biggest weaknesses, verbs and adpositions understanding, with more than 50% improvement on verbs and 46% improvement on adpositions.
To gain a better understanding of the model’s success and failures, we analyze the more frequent word failures and find some strong biases in vanilla CLIP. We show in Table IX some CLIP biases and how it change with 3VL. Note that removing less relevant tokens may enhance already existing biases. Yet, Table IX shows that this does not amplify the bias. Despite that it is clear that further reduction of the model bias is still required and our strategy just slighly improves it.
@@ V-G Ablation Study
We conduct an ablation study to assess the impact of different factors on the effectiveness of our proposed tree-based training technique. We examine the effects of different loss functions, different tree structures, and negatives generation. We examine the regularization effects of limiting our caption tree to a maximum depth or limiting the number of maximum negatives generated in each tree.
Different loss functions ablation. In Table III we show the impact of different loss functions on the performance of our model. Note that the addition of LoRA degrades a bit the performance on VL-checklist but reduces the forgetting effect (exhibited in the performance on ELEVATER [102]). Note that the addition of tree loss leads to a significant improvement over just using contrastive learning.
TABLE IX: CLIP fail rate on some pairs of positive-negative words on COCO test set.
@@
      “LLM prompt” - we generate co-hyponyms, antonyms, and random adjectives or adpositions with a prompt to FLAN-T5 (e.g., “find an opposite for the word: <>absent<>< >).
@@
  • Negatives from intermediate tree levels. We explored replacing also negatives from intermediate tree levels. For the example caption “several people standing in a green field together while flying kites” we get the following tree levels:
    1. 1.
      “several people”, “several animals”, “one people”
    2. 2.
      “several people and a green field”, “several animals and a green field”, “one people and a green field”, “several people and a blue field”, “several people and a green forest”
    3. 3.
      “several people standing in a green field”, “several animals standing in a green field”, “one people standing in a green field”, “several people standing in a blue field”, “several people standing in a green forest”, “several people standing out a green field”, “several people sitting in a green field”
    4. 4.
      “several people standing in a green field together while flying kites”, “several animals standing in a green field together while flying kites”, “one people standing in a green field together while flying kites”, “several people standing in a blue field together while flying kites”, “several people standing in a green forest together while flying kites”, “several people standing out a green field together while flying kites”, “several people sitting in a green field together while flying kites”, “several people standing in a green field together while soaring kites”, “several people standing in a green field together while flying sales”
@@
In Table X we show Top-1 accuracy results on VL-Checklist with six such tree variants. (i) “Basic WN” - “Basic” tree structure with WordNet negatives generation, (ii) “Incremental WN” - “Incremental” tree structure with WordNet negatives generation, (iii) “WN+LLM prompt+mask” - “Incremental” tree structure with “LLM prompt” for opposites generation, T5 mask completion for words with no opposite and WordNet replacement if T5 mask completion generates a word with similar meaning. (iv) “WN+LLM prompt+inter negs” - “Incremental” tree structure with “LLM prompt” for opposites generation and WordNet replacement if no opposite exists while using Negatives from intermediate tree levels. (v) “WN+LLM prompt+mix inter negs” - “Incremental” tree structure with “LLM prompt” for opposites generation and WordNet replacement if no opposite exists and mixing trees with Negatives from intermediate tree levels and without. (vi) “WN+LLM prompt (3VL)” - “Incremental” tree structure with “LLM prompt” for opposites generation and WordNet replacement if no opposite exists (without Negatives from intermediate tree levels).
@@
For each such variant, we report the highest result attained by hyperparameter sweep on the validation set of MSCOCO. We found that LLM mask completion oftentimes generated positive caption alternatives instead of negative captions and we did not use it in the final model. Best results were achieved using (vi) “WN+LLM prompt” and therefore, we used it in the final 3VL model.
@@
TABLE X: Top-1 accuracy on VL-Checklist for different caption trees. (i) “Basic WN” - Basic tree structure with WordNet replacement, (ii) “Incremental WN” - Incremental tree structure with WordNet replacement, (iii) “WN+LLM prompt+mask” - Incremental tree, FLAN-T5 prompt for opposites, T5 mask completion if opposite doesn’t exist and WordNet replacement if mask completion did not generate a word with a different meaning, (iv) “WN+LLM prompt+inter negs” - Incremental tree, FLAN-T5 prompt for opposites, and WordNet replacement if the opposite does not exist and using Negatives from intermediate tree levels, (v) “WN+LLM prompt+mix inter negs” - Incremental tree, FLAN-T5 prompt for opposites, and WordNet replacement if the opposite does not exist and mixing trees with Negatives from intermediate tree levels and without, (vi) “WN+LLM prompt (3VL)” - Incremental tree, FLAN-T5 prompt for opposites, and WordNet replacement if the opposite does not exist (without Negatives from intermediate tree levels).
@@
Tree Method Att color Att material Att size Att action Att state Rel action Rel spatial Avg Basic WN 70.54 75.79 69.72 75.95 70.46 81.81 73.94 74.03 Incremental WN 74.41 79.23 70.13 78.12 71.63 83.81 75.70 76.15 WN+LLM prompt+mask 72.31 77.53 69.98 76.69 69.39 83.43 76.87 75.17 WN+LLM prompt+inter negs 73.91 77.81 68.94 76.11 71.94 80.59 78.18 75.35 WN+LLM prompt+mix inter negs 73.79 78.10 69.05 76.64 72.58 80.67 78.24 5.58 WN+LLM prompt (3VL) 75.29 82.63 70.90 81.10 75.03 81.69 81.14 78.25
@@ V-H Maximum tree depth ablation
In Table XI we report Top-1 accuracy on the VL-Checklist when constraining the 3VL tree to maximum depth. We note that we don’t limit here the number of generated negatives per caption, just the structure of the tree. So, for example, a tree of depth 1 will contain the same number of negatives as in the full caption tree, just in a single level. We can see that even though the number of negatives remains the same, the increase in tree depth contributes to improved performance.